OpenStudy (anonymous):

Simplify the given expression. √-6(2+√-8) ?

OpenStudy (unklerhaukus):

√-6(2+√-8) =√(-12-12i√2) =√(-12(1-i√2) =2√(3(i√2-1)) =..

OpenStudy (anonymous):

2√6

OpenStudy (anonymous):

?

OpenStudy (unklerhaukus):

i got stuck